Can anyone explain to me how to get the manufacturing overhead number?  

Thank you!

Trees Sod
Selling Price Per Unit $                     12.99 $                                     0.38
Direct Materials Per Unit $                        0.99 $                                     0.01
Direct Labor Per Unit $                        2.00 $                                     0.09
Direct Labor Hours Per Unit 0.2 0.009
Estimated Annual Production and Sales 8,908 1,556,851 units
Estimated Overhead (last year's numbers)
Planting 47,000
Irrigation 19,000
Product Support 19,046
Harvesting 21,567
Inspections 4,564
Property Taxes 6,754
TOTAL 117,931
Estimated Total Direct Labor Hours 15793.259 DO NOT ROUND
Predetermined Overhead Rate 7.467173178 DO NOT ROUND
Traditional Costing
Sales $       115,714.92 $                     591,603.38
Direct Materials $             8,818.92 $                        15,568.51
Direct Labor $          17,816.00 $                     140,116.59
Manufacturing Overhead $          13,303.52                          104,627.48 HOW??
Product Margin $          75,776.48 $                     331,290.80
per unit $                        8.51 $                                     0.21
Answer #1

rate positively ..

Computation of overhead cost for department
Total cost
Planting 47,000
Irrigation 19,000
Product Support 19,046
Harvesting 21,567
Inspections 4,564
Property Taxes 6,754
TOTAL 117,931
Total hour
Trees =0.2*8908     1,781.60
Sod =1,556,851*.009 14,011.66
Total hour 15,793.26
Overhead rate = 117931/15793.26         7.467
therefore overhead = Overhead rate * number of hr
Trees overhead = 7.467*1781.6 13303.52
Sod overhead =7.467*14011.66 104627.48